How do we simplify 8 to thee 1/3 power? This is the problem I am currently working on so if you could give me a different example that is explained and answered I would be pleased! thankyou
Answer by vasu2qute(17)

Solution:
8^(1/3) =
(2^3)^(1/3)= 2^(3x1/3)=2^1=2